# Advancing Adversarial Suffix Transfer Learning on Aligned Large Language Models

### v1

@@ Abstract
Language Language Models (LLMs) face safety concerns due to potential misuse by malicious users. Recent red-teaming efforts have identified adversarial suffixes capable of jailbreaking LLMs using the gradient-based search algorithm Greedy Coordinate Gradient (GCG). However, GCG struggles with computational inefficiency, limiting further investigations regarding suffix transferability and scalability across models and data. In this work, we bridge the connection between search efficiency and suffix transferability. We propose a two-stage transfer learning framework, DeGCG, which decouples the search process into behavior-agnostic pre-searching and behavior-relevant post-searching. Specifically, we employ direct first target token optimization in pre-searching to facilitate the search process. We apply our approach to cross-model, cross-data, and self-transfer scenarios. Furthermore, we introduce an interleaved variant of our approach, i-DeGCG, which iteratively leverages self-transferability to accelerate the search process. Experiments on HarmBench demonstrate the efficiency of our approach across various models and domains. Notably, our i-DeGCG outperforms the baseline on Llama2-chat-7b with ASRs of $43.9$ ($+22.2$) and $39.0$ ($+19.5$) on valid and test sets, respectively. Further analysis on cross-model transfer indicates the pivotal role of first target token optimization in leveraging suffix transferability for efficient searching.
@@ 2.2 Jailbreak Attacks on Aligned LLMs
Existing red teaming has dedicated substantial efforts to identifying various jailbreak attacks. Initial jailbreak attacks involve the manual crafting of input prompts. A notable instance is the “Do-Anything-Now” attack, which is implemented by compelling LLMs to play a role that can do anything and respond to any query without refusal, thus bypassing safety constraints (Albert, 2023; Liu et al., 2023b). Subsequent advancements have automated the creation of these stealthy prompts (Liu et al., 2023a; Zhu et al., 2023). Additionally, adversarial prompts have been identified in GCG, which utilizes gradient information to automatically generate effective adversarial prompts (Zou et al., 2023; Shin et al., 2020). Furthermore, their results indicate the transferability and universality of these adversarial prompts. Recent work has also unveiled jailbreak attacks within the context of multilingual scenarios (Deng et al., 2023) and non-natural languages such as ciphers (Yuan et al., 2023), highlighting the risk for all open-source LLMs with modified decoding strategies (Huang et al., 2023). Our work focuses on adversarial suffix transferring learning across aligned LLMs and associates transferability with search efficiency.

### v2

@@ Abstract
Language Language Models (LLMs) face safety concerns due to potential misuse by malicious users. Recent red-teaming efforts have identified adversarial suffixes capable of jailbreaking LLMs using the gradient-based search algorithm Greedy Coordinate Gradient (GCG). However, GCG struggles with computational inefficiency, limiting further investigations regarding suffix transferability and scalability across models and data. In this work, we bridge the connection between search efficiency and suffix transferability. We propose a two-stage transfer learning framework, DeGCG, which decouples the search process into behavior-agnostic pre-searching and behavior-relevant post-searching. Specifically, we employ direct first target token optimization in pre-searching to facilitate the search process. We apply our approach to cross-model, cross-data, and self-transfer scenarios. Furthermore, we introduce an interleaved variant of our approach, i-DeGCG, which iteratively leverages self-transferability to accelerate the search process. Experiments on HarmBench demonstrate the efficiency of our approach across various models and domains. Notably, our i-DeGCG outperforms the baseline on Llama2-chat-7b with ASRs of $43.9$ ($+22.2$) and $39.0$ ($+19.5$) on valid and test sets, respectively. Further analysis on cross-model transfer indicates the pivotal role of first target token optimization in leveraging suffix transferability for efficient searching [^1] Code is publicly available at https://github.com/Waffle-Liu/DeGCG.
@@ 2.2 Jailbreak Attacks on Aligned LLMs
Existing red teaming has dedicated substantial efforts to identifying various jailbreak attacks. Initial jailbreak attacks involve the manual crafting of input prompts. A notable instance is the “Do-Anything-Now” attack, which is implemented by compelling LLMs to play a role that can do anything and respond to any query without refusal, thus bypassing safety constraints (Albert, 2023; Liu et al., 2023b). Subsequent advancements have automated the creation of these stealthy prompts with controllability (Liu et al., 2023a; Zhu et al., 2023; Guo et al., 2024). Additionally, adversarial prompts have been identified in GCG, which utilizes gradient information to automatically generate effective adversarial prompts (Zou et al., 2023; Shin et al., 2020; Zhao et al., 2024). Furthermore, their results indicate the transferability and universality of these adversarial prompts. Recent work has also unveiled jailbreak attacks within the context of multilingual scenarios (Deng et al., 2023), and non-natural languages such as ciphers (Yuan et al., 2023), in-context learning (Qiang et al., 2023), highlighting the risk for all open-source LLMs with modified decoding strategies (Huang et al., 2023). Our work focuses on adversarial suffix transferring learning across aligned LLMs and associates transferability with search efficiency.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank anonymous reviewers for their valuable suggestions. The authors also thank William Yang Wang, Xianjun Yang, Yiran Zhao, Keyu Duan, and Leon Lin for their helpful discussions. The computational work for this article was partially performed on resources of the National Supercomputing Centre (NSCC), Singapore [^2] https://www.nscc.sg/.
@@ A.2 Implementation Details
A.3 More results of Cross-Data Transferring
We provide more results of cross-data transferring on OpenChat-3.5-7b in Fig. 5 to supplement our conclusion discussed in Sec. 4.2.
Figure 5: Performance comparison (ASR) in Cross-Data Transferring across different behavior types in HarmBench. We report the results of OpenChat-3.5-7b on both the Validation and the Test sets.